# Deep End-to-end Adaptive k-Space Sampling, Reconstruction, and Registration for Dynamic MRI

### v1

@@ 0.B.1 Datasets
For external validation, the aorta dataset from the CMRxRecon 2024 challenge [50] was employed. This dataset also contains fully sampled, ECG-triggered multi-coil acquisitions segmented into 12 temporal frames. Transverse and sagittal views of the aorta were provided, with the subset used for inference consisting of 111 scans and 883 dynamic slices.
Figure 11 illustrates representative fully sampled reconstructed images from each dataset, showcasing the diversity of anatomical views.
@@
Figure 11: Representative fully sampled images across cardiac phases and views from each dataset.
0.B.2 Undersampling
0.B.2.1 Learned Sampling
@@ 0.B.2.2 Fixed Sampling
Figure 12: Representative examples of (fixed) sampling patterns.
@@ 0.B.3 Evaluation Metrics
0.B.4 Data Preprocessing and Postprocessing
0.B.4.1 Zero-padding
@@ 0.B.4.2 Normalization
$s=\text{quantile}_{99.5}\Big{(}\text{flatten}(|\mathbf{y}_{\text{mov}}^{\mathbf{M}^{\text{acs}}}|)\Big{)}.$
0.B.4.3 Postprocessing Cropping
@@
0.B.5 Hyperparameter Choices
0.B.5.1 Warping Transform
@@ 0.B.5.2 Registration Comparisons
    Our proposed registration model using a larger U-Net with enhanced capacity, featuring four scales with 32, 64, 128, and 252 filters at each scale.
@@
Appendix 0.C Additional Results
0.C.1 Additional Tables
@@
Reconstruction Model Sampling Type Sampling Initialization Registration Module Joint Loss Loss Weights 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 ×
@@
$\alpha$ $\beta$ SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
vSHARP Adaptive Learned 1 1 0.909 $\pm$ 0.041 33.51 $\pm$ 3.19 0.025 $\pm$ 0.015 0.910 $\pm$ 0.041 33.63 $\pm$ 3.18 0.024 $\pm$ 0.014 0.909 $\pm$ 0.042 33.67 $\pm$ 3.16 0.024 $\pm$ 0.014
vSHARP Adaptive Learned (L) 1 1 0.908 $\pm$ 0.041 33.43 $\pm$ 3.14 0.025 $\pm$ 0.015 0.909 $\pm$ 0.041 33.58 $\pm$ 3.13 0.024 $\pm$ 0.014 0.909 $\pm$ 0.041 33.64 $\pm$ 3.12 0.024 $\pm$ 0.014
vSHARP Adaptive Voxelmorph 1 1 0.889 $\pm$ 0.046 32.16 $\pm$ 2.92 0.033 $\pm$ 0.017 0.886 $\pm$ 0.047 32.13 $\pm$ 2.91 0.033 $\pm$ 0.017 0.882 $\pm$ 0.048 32.04 $\pm$ 2.89 0.034 $\pm$ 0.017
vSHARP Adaptive Transfmorph 1 1 0.886 $\pm$ 0.046 31.37 $\pm$ 3.02 0.040 $\pm$ 0.022 0.886 $\pm$ 0.046 31.47 $\pm$ 3.00 0.039 $\pm$ 0.022 0.886 $\pm$ 0.046 31.50 $\pm$ 2.95 0.039 $\pm$ 0.021
vSHARP Adaptive Learned 1 3 0.939 $\pm$ 0.031 36.99 $\pm$ 3.26 0.011 $\pm$ 0.007 0.936 $\pm$ 0.032 36.79 $\pm$ 3.22 0.012 $\pm$ 0.008 0.932 $\pm$ 0.034 36.49 $\pm$ 3.19 0.013 $\pm$ 0.008
vSHARP Adaptive Learned 1 1 0.908 $\pm$ 0.041 33.48 $\pm$ 3.14 0.025 $\pm$ 0.014 0.908 $\pm$ 0.042 33.56 $\pm$ 3.14 0.025 $\pm$ 0.014 0.907 $\pm$ 0.042 33.56 $\pm$ 3.12 0.025 $\pm$ 0.014
vSHARP Adaptive Learned 1 1 0.907 $\pm$ 0.042 $3$$\pm$ 3.15 0.026 $\pm$ 0.015 0.907 $\pm$ 0.042 33.45 $\pm$ 3.15 0.025 $\pm$ 0.015 0.906 $\pm$ 0.042 33.47 $\pm$ 3.12 0.025 $\pm$ 0.015
vSHARP Adaptive Learned 1 1 0.905 $\pm$ 0.043 32.95 $\pm$ 3.15 0.028 $\pm$ 0.017 0.906 $\pm$ 0.043 33.08 $\pm$ 3.15 0.028 $\pm$ 0.016 0.904 $\pm$ 0.043 33.13 $\pm$ 3.12 0.027 $\pm$ 0.016
vSHARP Equi NA Learned 1 1 0.897 $\pm$ 0.045 32.47 $\pm$ 3.08 0.031 $\pm$ 0.018 0.898 $\pm$ 0.045 32.65 $\pm$ 3.08 0.030 $\pm$ 0.017 0.893 $\pm$ 0.046 32.61 $\pm$ 3.00 0.030 $\pm$ 0.017
vSHARP Equi NA Learned 1 1 0.896 $\pm$ 0.045 32.31 $\pm$ 3.07 0.033 $\pm$ 0.019 0.895 $\pm$ 0.046 32.44 $\pm$ 3.07 0.032 $\pm$ 0.018 0.890 $\pm$ 0.047 32.38 $\pm$ 2.99 0.032 $\pm$ 0.018
vSHARP $k$t-equi NA Learned 1 1 0.899 $\pm$ 0.045 32.59 $\pm$ 3.14 0.031 $\pm$ 0.018 0.900 $\pm$ 0.045 32.77 $\pm$ 3.14 0.030 $\pm$ 0.017 0.893 $\pm$ 0.047 32.73 $\pm$ 3.09 0.030 $\pm$ 0.017
vSHARP $k$t-equi NA Learned 1 1 0.895 $\pm$ 0.046 32.24 $\pm$ 3.10 0.033 $\pm$ 0.019 0.896 $\pm$ 0.046 32.40 $\pm$ 3.09 0.032 $\pm$ 0.018 0.890 $\pm$ 0.048 32.37 $\pm$ 3.05 0.032 $\pm$ 0.018
vSHARP Optimized Learned 1 1 0.912 $\pm$ 0.040 33.76 $\pm$ 3.14 0.023 $\pm$ 0.013 0.912 $\pm$ 0.040 33.87 $\pm$ 3.14 0.023 $\pm$ 0.013 0.910 $\pm$ 0.041 33.87 $\pm$ 3.11 0.023 $\pm$ 0.013
VarNet Adaptive Learned 1 1 0.886 $\pm$ 0.052 33.15 $\pm$ 2.97 0.026 $\pm$ 0.013 0.870 $\pm$ 0.056 32.44 $\pm$ 2.87 0.030 $\pm$ 0.015 0.863 $\pm$ 0.055 32.15 $\pm$ 2.73 0.032 $\pm$ 0.015
vSHARP Adaptive OFILK NA 1 NA 0.890 $\pm$ 0.044 32.20 $\pm$ 2.89 0.033 $\pm$ 0.018 0.889 $\pm$ 0.044 32.21 $\pm$ 2.89 0.033 $\pm$ 0.018 0.887 $\pm$ 0.045 32.17 $\pm$ 2.87 0.033 $\pm$ 0.018
vSHARP Adaptive OFTVL1 NA 1 NA 0.893 $\pm$ 0.042 30.69 $\pm$ 3.08 0.049 $\pm$ 0.030 0.892 $\pm$ 0.042 30.66 $\pm$ 3.06 0.049 $\pm$ 0.030 0.891 $\pm$ 0.042 30.63 $\pm$ 3.05 0.049 $\pm$ 0.030
vSHARP Adaptive DEMONS NA 1 NA 0.892 $\pm$ 0.044 30.77 $\pm$ 2.87 0.045 $\pm$ 0.023 0.891 $\pm$ 0.045 30.80 $\pm$ 2.86 0.045 $\pm$ 0.023 0.890 $\pm$ 0.045 30.82 $\pm$ 2.85 0.045 $\pm$ 0.023
vSHARP Equi NA OFILK NA 1 NA 0.888 $\pm$ 0.045 32.14 $\pm$ 2.89 0.033 $\pm$ 0.018 0.886 $\pm$ 0.045 32.14 $\pm$ 2.88 0.033 $\pm$ 0.018 0.880 $\pm$ 0.048 32.01 $\pm$ 2.84 0.034 $\pm$ 0.018
vSHARP Equi NA OFTVL1 NA 1 NA 0.891 $\pm$ 0.042 30.58 $\pm$ 3.09 0.050 $\pm$ 0.031 0.889 $\pm$ 0.043 30.55 $\pm$ 3.07 0.050 $\pm$ 0.031 0.883 $\pm$ 0.045 30.37 $\pm$ 3.01 0.052 $\pm$ 0.032
vSHARP Equi NA DEMONS NA 1 NA 0.890 $\pm$ 0.045 30.74 $\pm$ 2.88 0.046 $\pm$ 0.023 0.889 $\pm$ 0.046 30.78 $\pm$ 2.87 0.045 $\pm$ 0.023 0.884 $\pm$ 0.047 30.79 $\pm$ 2.85 0.045 $\pm$ 0.023
vSHARP $k$t-equi NA OFILK NA 1 NA 0.889 $\pm$ 0.045 32.18 $\pm$ 2.91 0.033 $\pm$ 0.018 0.886 $\pm$ 0.046 32.16 $\pm$ 2.91 0.033 $\pm$ 0.018 0.879 $\pm$ 0.049 32.03 $\pm$ 2.88 0.034 $\pm$ 0.018
vSHARP $k$t-equi NA OFTVL1 NA 1 NA 0.891 $\pm$ 0.042 30.62 $\pm$ 3.09 0.049 $\pm$ 0.030 0.888 $\pm$ 0.044 30.51 $\pm$ 3.08 0.051 $\pm$ 0.031 0.881 $\pm$ 0.046 30.33 $\pm$ 3.04 0.053 $\pm$ 0.032
vSHARP $k$t-equi NA DEMONS NA 1 NA 0.891 $\pm$ 0.045 30.77 $\pm$ 2.89 0.045 $\pm$ 0.023 0.889 $\pm$ 0.046 30.81 $\pm$ 2.90 0.045 $\pm$ 0.023 0.884 $\pm$ 0.048 30.84 $\pm$ 2.87 0.045 $\pm$ 0.023
@@
Table 1: Quantitative results for various configurations on the cardiac cine test set (under phase-specific settings).
@@
Reconstruction Model Sampling Type Sampling Initialization Registration Module Joint Loss Loss Weights 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 ×

### v2

@@ B.1 Datasets
For external validation, the aorta dataset from the CMRxRecon 2024 challenge [51, 50] was employed. This dataset also contains fully sampled, ECG-triggered multi-coil acquisitions segmented into 12 temporal frames. Transverse and sagittal views of the aorta were provided, with the subset used for inference consisting of 111 scans and 883 dynamic slices.
Figure B1 illustrates representative fully sampled reconstructed images from each dataset, showcasing the diversity of anatomical views.
@@
Figure B1: Representative fully sampled images across cardiac phases and views from each dataset.
B.2 Undersampling
B.2.1 Learned Sampling
@@ B.2.2 Fixed Sampling
Figure B2: Representative examples of (fixed) sampling patterns.
@@ B.3 Evaluation Metrics
B.4 Data Preprocessing and Postprocessing
B.4.1 Zero-padding
@@ B.4.2 Normalization
$s=\text{quantile}_{99.5}\Big{(}\text{flatten}(|\mathbf{y}_{\text{mov}}^{\mathbf{M}^{\text{acs}}}|)\Big{)}.$ (B4)
B.4.3 Postprocessing Cropping
@@
B.5 Hyperparameter Choices
B.5.1 Warping Transform
@@ B.5.2 Registration Comparisons
Appendix C Additional Results
C.1 Additional Experiments
Due to space constraints in the main paper, we present additional experiments here to further validate our proposed pipeline across various settings, analyzing different aspects of its performance. We report all average results, including those from the main paper, in Tables C1 to C8. These results include evaluations on both the cine test set—sharing the same data distribution as the training set—and an out-of-distribution aorta inference set.
First, we extend our registration component experiments by considering the methods described in the main paper while testing additional configurations. Specifically, we assess performance using non-adaptive sampling strategies, such as equispaced and $k$t-equispaced trajectories, alongside adaptive sampling with initialization. These experiments are conducted for both phase-specific and unified sampling approaches. Additionally, we expand our reconstruction model comparisons to include the unified sampling setting, complementing the phase-specific evaluations performed in the original experiments.
Furthermore, we investigate the limiting case where $\alpha=0$, effectively setting the reconstruction loss weight in the joint loss calculation to zero, thereby omitting reconstruction loss optimization. We also provide qualitative visualizations of this scenario in Figures C7 and C8.
Finally, to analyze the impact of the smoothness/regularization component $\mathcal{L}_{\text{smooth}}$ within the registration loss, we conduct phase-specific experiments by removing this term. We also explore different combinations of $\alpha$ and $\beta$ in the absence of this component to assess its contribution to the overall performance.
C.2 Additional Tables
@@
Registration Model Reconstruction Model Sampling Type Sampling Initialization Loss Details 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × Inference Time
@@
$\mathcal{L}_{\mathrm{sim}},\mathcal{L}_{\mathrm{reg}}$ $\alpha$ $\beta$ $\mathcal{L}_{\mathrm{smooth}}$ E2E Loss SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NMSE ($\downarrow$)
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.909\pm 0.041$ $33.51\pm 3.19$ $0.025\pm 0.015$ $0.910\pm 0.041$ $33.63\pm 3.18$ $0.024\pm 0.014$ $0.909\pm 0.042$ $33.67\pm 3.16$ $0.024\pm 0.014$ 10.16
Voxelmorph vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.889\pm 0.046$ $32.16\pm 2.92$ $0.033\pm 0.017$ $0.886\pm 0.047$ $32.13\pm 2.91$ $0.033\pm 0.017$ $0.882\pm 0.048$ $32.04\pm 2.89$ $0.034\pm 0.017$ 10.74
Transfmorph vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.886\pm 0.046$ $31.37\pm 3.02$ $0.040\pm 0.022$ $0.886\pm 0.046$ $31.47\pm 3.00$ $0.039\pm 0.022$ $0.886\pm 0.046$ $31.50\pm 2.95$ $0.039\pm 0.021$ 10.11
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 3.0 $0.939\pm 0.031$ $36.99\pm 3.26$ $0.011\pm 0.007$ $0.936\pm 0.032$ $36.79\pm 3.22$ $0.012\pm 0.008$ $0.932\pm 0.034$ $36.49\pm 3.19$ $0.013\pm 0.008$ 10.90
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 2.0 $0.931\pm 0.033$ $35.76\pm 3.17$ $0.015\pm 0.009$ $0.930\pm 0.034$ $35.74\pm 3.16$ $0.015\pm 0.009$ $0.927\pm 0.034$ $35.60\pm 3.12$ $0.015\pm 0.009$ 10.40
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.5 2.0 $0.918\pm 0.038$ $34.23\pm 3.14$ $0.021\pm 0.012$ $0.918\pm 0.038$ $34.33\pm 3.12$ $0.021\pm 0.012$ $0.916\pm 0.038$ $34.33\pm 3.10$ $0.021\pm 0.012$ 11.78
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 0.0 1.0 $0.987\pm 0.006$ $44.15\pm 2.53$ $0.002\pm 0.001$ $0.985\pm 0.007$ $43.78\pm 2.45$ $0.002\pm 0.001$ $0.984\pm 0.008$ $43.22\pm 2.41$ $0.003\pm 0.001$ 10.69
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 2.0 1.0 $0.893\pm 0.046$ $32.13\pm 3.07$ $0.034\pm 0.019$ $0.893\pm 0.046$ $32.19\pm 3.07$ $0.033\pm 0.019$ $0.893\pm 0.047$ $32.22\pm 3.06$ $0.033\pm 0.019$ 10.87
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.881\pm 0.051$ $31.47\pm 3.06$ $0.039\pm 0.022$ $0.881\pm 0.051$ $31.50\pm 3.06$ $0.039\pm 0.022$ $0.881\pm 0.051$ $31.63\pm 3.04$ $0.038\pm 0.021$ 10.89
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.925\pm 0.033$ $33.66\pm 3.05$ $0.024\pm 0.014$ $0.926\pm 0.033$ $33.72\pm 3.04$ $0.024\pm 0.014$ $0.927\pm 0.032$ $33.82\pm 3.04$ $0.023\pm 0.014$ 10.99
Learned vSHARP Adaptive $\mathcal{L}_{1}$ 1.0 1.0 $0.886\pm 0.049$ $31.69\pm 3.10$ $0.038\pm 0.021$ $0.887\pm 0.048$ $31.80\pm 3.09$ $0.037\pm 0.021$ $0.886\pm 0.049$ $31.85\pm 3.08$ $0.036\pm 0.020$ 11.91
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim2D}}$ 1.0 1.0 $0.896\pm 0.043$ $31.75\pm 2.92$ $0.037\pm 0.020$ $0.895\pm 0.043$ $31.83\pm 2.90$ $0.036\pm 0.019$ $0.893\pm 0.044$ $31.81\pm 2.88$ $0.036\pm 0.019$ 10.47
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.948\pm 0.024$ $35.54\pm 3.18$ $0.016\pm 0.010$ $0.949\pm 0.023$ $35.74\pm 3.18$ $0.015\pm 0.009$ $0.950\pm 0.023$ $35.88\pm 3.17$ $0.015\pm 0.009$ 10.66
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 2.0 $0.967\pm 0.015$ $38.02\pm 3.23$ $0.009\pm 0.005$ $0.967\pm 0.016$ $38.08\pm 3.21$ $0.009\pm 0.005$ $0.967\pm 0.016$ $38.08\pm 3.19$ $0.009\pm 0.005$ 10.40
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 2.0 1.0 $0.933\pm 0.030$ $34.18\pm 3.11$ $0.022\pm 0.013$ $0.935\pm 0.029$ $34.34\pm 3.10$ $0.021\pm 0.013$ $0.937\pm 0.028$ $34.47\pm 3.10$ $0.020\pm 0.013$ 10.95
Learned vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.907\pm 0.042$ $33.33\pm 3.15$ $0.026\pm 0.015$ $0.907\pm 0.042$ $33.45\pm 3.15$ $0.025\pm 0.015$ $0.906\pm 0.042$ $33.47\pm 3.12$ $0.025\pm 0.015$ 10.18
Learned vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.897\pm 0.045$ $32.47\pm 3.08$ $0.031\pm 0.018$ $0.898\pm 0.045$ $32.65\pm 3.08$ $0.030\pm 0.017$ $0.893\pm 0.046$ $32.61\pm 3.00$ $0.030\pm 0.017$ 9.93
Learned vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.899\pm 0.045$ $32.59\pm 3.14$ $0.031\pm 0.018$ $0.900\pm 0.045$ $32.77\pm 3.14$ $0.030\pm 0.017$ $0.893\pm 0.047$ $32.73\pm 3.09$ $0.030\pm 0.017$ 9.91
Learned vSHARP Optimized $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.912\pm 0.040$ $33.76\pm 3.14$ $0.023\pm 0.013$ $0.912\pm 0.040$ $33.87\pm 3.14$ $0.023\pm 0.013$ $0.910\pm 0.041$ $33.87\pm 3.11$ $0.023\pm 0.013$ 10.85
Learned VarNet Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 $0.886\pm 0.052$ $33.15\pm 2.97$ $0.026\pm 0.013$ $0.870\pm 0.056$ $32.44\pm 2.87$ $0.030\pm 0.015$ $0.863\pm 0.055$ $32.15\pm 2.73$ $0.032\pm 0.015$ 10.59
OptFlowILK vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.890\pm 0.044$ $32.20\pm 2.89$ $0.033\pm 0.018$ $0.889\pm 0.044$ $32.21\pm 2.89$ $0.033\pm 0.018$ $0.887\pm 0.045$ $32.17\pm 2.87$ $0.033\pm 0.018$ 27.27
OptFlowTVL1 vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.893\pm 0.042$ $30.69\pm 3.08$ $0.049\pm 0.030$ $0.892\pm 0.042$ $30.66\pm 3.06$ $0.049\pm 0.030$ $0.891\pm 0.042$ $30.63\pm 3.05$ $0.049\pm 0.030$ 30.20
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.892\pm 0.044$ $30.77\pm 2.87$ $0.045\pm 0.023$ $0.891\pm 0.045$ $30.80\pm 2.86$ $0.045\pm 0.023$ $0.890\pm 0.045$ $30.82\pm 2.85$ $0.045\pm 0.023$ 29.31
OptFlowILK vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.891\pm 0.044$ $32.19\pm 2.90$ $0.033\pm 0.018$ $0.890\pm 0.044$ $32.21\pm 2.89$ $0.033\pm 0.018$ $0.888\pm 0.045$ $32.19\pm 2.88$ $0.033\pm 0.018$ 27.18
OptFlowTVL1 vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.895\pm 0.041$ $30.79\pm 3.12$ $0.048\pm 0.030$ $0.894\pm 0.041$ $30.76\pm 3.12$ $0.048\pm 0.030$ $0.892\pm 0.042$ $30.72\pm 3.11$ $0.048\pm 0.030$ 30.16
DEMONS vSHARP Adaptive $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.892\pm 0.044$ $30.75\pm 2.89$ $0.046\pm 0.023$ $0.892\pm 0.045$ $30.80\pm 2.88$ $0.045\pm 0.023$ $0.891\pm 0.045$ $30.83\pm 2.87$ $0.045\pm 0.023$ 28.92
OptFlowILK vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.888\pm 0.045$ $32.14\pm 2.89$ $0.033\pm 0.018$ $0.886\pm 0.045$ $32.14\pm 2.88$ $0.033\pm 0.018$ $0.880\pm 0.048$ $32.01\pm 2.84$ $0.034\pm 0.018$ 26.58
OptFlowTVL1 vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.891\pm 0.042$ $30.58\pm 3.09$ $0.050\pm 0.031$ $0.889\pm 0.043$ $30.55\pm 3.07$ $0.050\pm 0.031$ $0.883\pm 0.045$ $30.37\pm 3.01$ $0.052\pm 0.032$ 30.08
DEMONS vSHARP Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.890\pm 0.045$ $30.74\pm 2.88$ $0.046\pm 0.023$ $0.889\pm 0.046$ $30.78\pm 2.87$ $0.045\pm 0.023$ $0.884\pm 0.047$ $30.79\pm 2.85$ $0.045\pm 0.023$ 29.00
OptFlowILK vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.889\pm 0.045$ $32.18\pm 2.91$ $0.033\pm 0.018$ $0.886\pm 0.046$ $32.16\pm 2.91$ $0.033\pm 0.018$ $0.879\pm 0.049$ $32.03\pm 2.88$ $0.034\pm 0.018$ 26.46
OptFlowTVL1 vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.891\pm 0.042$ $30.62\pm 3.09$ $0.049\pm 0.030$ $0.888\pm 0.044$ $30.51\pm 3.08$ $0.051\pm 0.031$ $0.881\pm 0.046$ $30.33\pm 3.04$ $0.053\pm 0.032$ 29.74
DEMONS vSHARP kt-Equispaced NA $\mathcal{L}_{\mathrm{ssim}}+\mathcal{L}_{1}$ 1.0 1.0 NA $0.891\pm 0.045$ $30.77\pm 2.89$ $0.045\pm 0.023$ $0.889\pm 0.046$ $30.81\pm 2.90$ $0.045\pm 0.023$ $0.884\pm 0.048$ $30.84\pm 2.87$ $0.045\pm 0.023$ 28.78
@@
Table C1: Registration quantitative results for various configurations on the cardiac cine test set (under phase-specific settings).
@@
Registration Model Reconstruction Model Sampling Type Sampling Initialization Loss Details 4×4$\times$4 × 6×6$\times$6 × 8×8$\times$8 × Inference Time